# CoTAL: Human-in-the-Loop Prompt Engineering, Chain-of-Thought Reasoning, and Active Learning for Generalizable Formative Assessment Scoring

### v2

@@ VI Analyzing RQ2: Educator and Student Feedback
RQ2 asked: What do teacher and student input reveal about the effectiveness, actionability, and impact of CoTAL’s formative feedback? We answer this question qualitatively by memoing key findings and using constant comparative analysis to analyze educators’ interviews and students’ surveys, respectively.
VI-A Educator Feedback
@@
We conducted semi-structured interviews with two classroom teachers, asking them to reflect on CoTAL’s scoring of several of their students’ Science Concepts and Reasoning responses. Both teachers reported that CoTAL achieved high scoring accuracy. They recognized the utility of LLMs in enhancing teaching efficiency and identifying student learning gaps, thereby guiding subsequent educational interventions. One educator personally undertook the Science Concepts and Reasoning assessment and received a score of 6 out of 9 from the LLM. She noted CoTAL’s effectiveness in pinpointing and explaining her mistakes and detecting her misunderstanding. The other educator emphasized CoTAL’s (and LLMs’, more generally) potential to reduce teacher bias by evaluating students based solely on their answers, not preconceptions about the student:
@@
This educator also underscored the value of collaboration between humans and artificial intelligence in education, particularly in teaching language models to recognize situations that necessitate teacher involvement:
@@
The two teachers proposed refinements to CoTAL, particularly suggesting that the LLM notify the teachers about students who need additional support and provide feedback that recommends subsequent actions to enhance their learning, such as study topics tailored to their knowledge gaps. One teacher outlined three key functions for an enhanced LLM grading system: (1) offering students constructive feedback for reflection, (2) sharing student performance data with educators, and (3) alerting teachers to notable insights in student submissions. This teacher emphasized the importance of the LLM asking thought-provoking questions to students to evaluate and enhance their conceptual understanding of science topic(s):
@@
…[the AI] doing things that make you more productive as the teacher because that’s the kind of stuff that you can do as a teacher, it just is so time intensive, times every kid…so taking some of that legwork out for the teacher…not replacing what the teacher does just, doing the legwork.
@@
VI-B Student Feedback
We conducted a focus group discussion and survey with a class of 23 students, asking whether they agreed with GPT-4’s scores and explanations using CoTAL to score their Science Concepts and Reasoning formative assessments. We also inquired about the usefulness of CoTAL’s feedback and their confidence in AI grading future assignments. The students were enthusiastic participants, providing a range of positive and negative insights into using LLMs as automated graders.
Overall, 61% of the students found CoTAL’s responses helpful, and 65% expressed confidence that an LLM would score future assignments well. Many students specifically noted that the LLM’s scores and explanations were helpful and accurate, feeling that the LLM adequately understood their responses:
@@ VI-B Student Feedback
I liked that it explained throughly [sic] what I did wrong and what I did right.
@@
The most frequent comment by students was that CoTAL’s describing how it awarded (or did not award) points and explaining their errors was helpful, particularly concerning the LLM citing evidence and tying it back to the rubric:
@@
…the chat does a relatively good job explaining what I did wrong I think I could work with the feedback to help me learn the content better.
@@
One way to mitigate this issue is to include examples of ideal responses alongside each scoring explanation so the students can compare their responses to the ground truth (as established by their teacher) and understand the differences in relation to the rubric. Students also suggested making the LLM’s tone less harsh and its responses shorter and less repetitive (“Be less harsh.”; “Stop repetition”; “I like how it explains all the things it had for requirements, but its a little long.”) Like their teachers, students expressed a desire for the LLM to acknowledge their achievements in addition to identifying areas of improvement (“Showing what the student did right”; “I think it can do better with showing what you did right.”).
@@
Regarding the effectiveness, actionability, and impact of CoTAL’s formative feedback, students focused on its ability to explain scoring decisions, which helped them identify the misunderstandings that resulted in incorrect responses. Understanding why the LLM awarded its scores allowed students to connect their response shortcomings to the rubric and address learning gaps. This “explainability” fosters student trust in AI systems, as prior work shows students are often reluctant to trust AI-generated grades without understanding the reasoning, hindering their willingness to act on LLM-generated feedback [43].
VI-C Answering RQ2
@@
Teacher interviews and student surveys revealed several insights into CoTAL’s efficacy and utility in classroom settings. Both groups acknowledged the LLM’s scoring accuracy and ability to explain why responses were correct or incorrect, identifying learning opportunities and interventions. They showed a willingness to accept LLMs in educational settings, particularly with human oversight, aligning with previous findings that ChatGPT enjoys a supportive attitude in academia [44, 45]. However, teachers and students noted the need for enhancements. 39% of the students did not find the feedback helpful, and 35% lacked confidence in the LLM’s grading capabilities. Some students wanted the LLM to pinpoint deficiencies and provide more detailed feedback on weaknesses. Teachers emphasized the need for more actionable feedback and “next steps” to address learning gaps and guide a deeper understanding of science topics.
@@
These insights suggest actionable steps to refine CoTAL. We plan to create prompts that highlight correct and incorrect parts of students’ answers. Recognizing correct elements of an answer fosters greater engagement and trust between students and LLMs. To help students grasp their mistakes, we will provide exemplary answers alongside LLM feedback, clarifying why they did not earn full points and which concepts were misunderstood. CoTAL will identify misunderstandings for deeper discussion and inform teachers so that they may support struggling students. We will enhance stakeholder involvement through participatory design during CoTAL Phase II (Response Scoring, Prompt Development, and Active Learning).
VII Discussion and Conclusions
Our work provides several implications for K-12 educational methodology and practice. Traditional automated grading methods, such as supervised fine-tuning, require large datasets and extensive GPU training. In contrast, prompt engineering techniques like WRVRT [33] do not take into account input from teachers or students and lack evaluation across multiple domains. CoTAL presents a significant advantage by aligning LLMs with teachers effectively and adapting responses to reflect the preferences of individual stakeholders in a generalizable way, utilizing only a small set of demonstration instances to incorporate human insight.
@@ VII Discussion and Conclusions
This is particularly valuable for subjective tasks such as feedback generation and formative assessment scoring, where teachers differ in their preferences, making a single fine-tuned LLM unlikely to align well with multiple educators. While much work has explored adapting LLMs for personalizing student interactions — typically through content and feedback generation tailored to individual learning styles [46] — CoTAL addresses a notable gap in learning sciences and technologies methodology by personalizing LLMs for educators.
@@
Teacher and student feedback revealed areas for improving stakeholder alignment and underscored the importance of fostering trust in AI systems before expecting real-world adoption. Students often prefer ChatGPT-assisted (i.e., human-in-the-loop) grading relative to fully automated methods [47, 48]. Jiang et al. (2024) [44] found that, while there is “general optimism” about AI’s potential to enhance teaching, concerns persist regarding ethics, discrimination, and regulatory gaps. Lee and Zhai (2024) [49] identified several “inappropriate” ChatGPT use cases where the LLM hallucinated internet material, further adding ChatGPT could be used in ways that “potentially undermine students’ practical inquiry skills over time.” This highlights the need for additional mechanisms that prioritize enhancing LLM reliability, trust, and curricular alignment—not just accuracy—which our findings reinforce.
@@
Hallucination reduction is essential. Huang et al.’s (2023) [42] survey on LLM hallucinations identified several hallucination causes, discussing reasoning failure where “[an LLM] may struggle to produce accurate results if multiple associations exist between questions,” even in instances where the LLM possesses the necessary knowledge. In our case, this manifested through context inconsistency errors due to the LLM conflating the absorption and absorption limit concepts in the Rules and Debugging Tasks. These concepts are inextricably linked and often appear in similar contexts, which challenges differentiation. Just as students struggled to distinguish between these concepts, we hypothesize that the LLM’s training data reflects similar ambiguities, which we will investigate in future work. Hallucination mitigation strategies such as retrieval-augmented generation (RAG) to ground responses in human-curated data and decoding strategies that prioritize adherence to facts and user instructions similarly warrant future investigation.
Our work is not without limitations. While we demonstrated CoTAL’s generalizability across science, computing, and engineering domains within an integrated STEM+C curriculum, further research is needed to evaluate its performance across additional domains and curricula. This study examines CoTAL’s scoring performance in a post hoc setting without assessing its real-time use in classrooms, limiting our ability to determine its impact on student learning. Additionally, human-in-the-loop prompt engineering can be time-consuming, raising concerns about CoTAL’s scalability and how much it reduces teachers’ workload in real-world classroom settings.
@@
Recently, we conducted a follow-up study in four classrooms, each with approximately 26 students, deploying CoTAL via a formative assessment agent. This agent enabled students to discuss their CoTAL-generated formative assessment scores, helping them understand their mistakes and how to improve their performance moving forward. While a comprehensive analysis of students’ learning gains and agent interactions is forthcoming, we observed substantial time savings with CoTAL compared to our earlier efforts involving human scoring, providing students feedback within hours instead of weeks.

### v3

@@ 6 Analyzing RQ2: Teacher and Student Feedback
RQ2 asked: What do teacher and student input reveal about the effectiveness, actionability, and impact of CoTAL’s formative feedback? We answer this question qualitatively by memoing key findings and using constant comparative analysis to analyze teachers’ interviews and students’ surveys, respectively.
6.1 Teacher Feedback
@@
We conducted semi-structured interviews with two classroom teachers, asking them to reflect on CoTAL’s scoring of several of their students’ Science Concepts and Reasoning responses. Both teachers reported that CoTAL achieved high scoring accuracy. They recognized the utility of LLMs in enhancing teaching efficiency and identifying student learning gaps, thereby guiding subsequent educational interventions. One teacher personally undertook the Science Concepts and Reasoning assessment and received a score of 6 out of 9 from the LLM. She noted CoTAL’s effectiveness in pinpointing and explaining her mistakes and detecting her misunderstanding. The other teacher emphasized CoTAL’s (and LLMs’, more generally) potential to reduce teacher bias by evaluating students based solely on their answers, not preconceptions about the student:
@@
This teacher also underscored the value of collaboration between humans and artificial intelligence in education, particularly in teaching language models to recognize situations that necessitate teacher involvement:
@@
The two teachers proposed refinements to CoTAL, particularly suggesting that the LLM notify the teachers about students who need additional support and provide feedback that recommends subsequent actions to enhance their learning, such as study topics tailored to their knowledge gaps. One teacher outlined three key functions for an enhanced LLM grading system: (1) offering students constructive feedback for reflection, (2) sharing student performance data with teachers, and (3) alerting teachers to notable insights in student submissions. This teacher emphasized the importance of the LLM asking thought-provoking questions to students to evaluate and enhance their conceptual understanding of science topic(s):
@@
…[the AI] doing things that make you more productive as the teacher because that’s the kind of stuff that you can do as a teacher, it just is so time intensive, times every kid…so taking some of that legwork out for the teacher…not replacing what the teacher does, just doing the legwork.
@@
6.2 Student Feedback
We conducted a focus group and survey with a class of 23 students to evaluate agreement with GPT-4’s scores and explanations generated using CoTAL for the Science Concepts and Reasoning formative assessments. We also solicited perceptions of the usefulness of CoTAL’s feedback and confidence in AI-based grading for future assignments. Students engaged enthusiastically and provided a range of constructive observations, including both positive and critical perspectives on employing LLMs as automated graders.
Overall, 61% of students reported that CoTAL’s responses were helpful, and 65% expressed confidence that an LLM would accurately grade future assignments. Many students specifically noted that the LLM’s scores and explanations were helpful and accurate, and that the model adequately understood their responses:
@@ 6.2 Student Feedback
I liked that it explained thoroughly [sic] what I did wrong and what I did right.
@@
The most frequent comment by students was that CoTAL’s description of how it awarded (or did not award) points and explaining their errors was helpful, particularly concerning the LLM citing evidence and tying it back to the rubric:
@@
…the chat does a relatively good job explaining what I did wrong, I think I could work with the feedback to help me learn the content better.
@@
One way to mitigate this issue is to include examples of ideal responses alongside each scoring explanation so the students can compare their responses to the ground truth (as established by their teacher) and understand the differences with the rubric. Students also suggested making the LLM’s tone less harsh and its responses shorter and less repetitive (“Be less harsh.”; “Stop repetition”; “I like how it explains all the things it had for requirements, but its a little long.”) Like their teachers, students expressed a desire for the LLM to acknowledge their achievements in addition to identifying areas of improvement (“Showing what the student did right”; “I think it can do better with showing what you did right.”).
@@
Regarding the effectiveness, actionability, and impact of CoTAL’s formative feedback, students emphasized its capacity to explain scoring decisions, which helped them pinpoint the misunderstandings underlying incorrect responses. By understanding why the LLM assigned particular scores, students were able to map deficiencies in their responses to specific rubric criteria and address corresponding learning gaps. This “explainability” appears to foster trust in AI systems; prior work indicates that students are often reluctant to accept AI-generated grades without transparent reasoning, which can reduce their willingness to act on LLM-generated feedback [59].
6.3 Answering RQ2
@@
Teacher interviews and student surveys yielded several insights into CoTAL’s efficacy and classroom utility. Both groups recognized the LLM’s scoring accuracy and its capacity to explain why responses were correct or incorrect, thereby identifying learning opportunities and potential interventions. Participants expressed willingness to adopt LLMs in educational contexts, particularly with human oversight, consistent with prior findings that ChatGPT enjoys a supportive attitude in academia [60, 61]. Nevertheless, areas for improvement were noted: 39% of students did not find the feedback helpful, and 35% lacked confidence in the LLM’s grading capabilities. Some students requested more detailed, diagnostic feedback that pinpointed deficiencies. At the same time, teachers emphasized the need for actionable guidance and “next steps” to address learning gaps and deepen understanding of science topics.
@@
These findings suggest concrete refinements to CoTAL. We plan to develop prompts that explicitly highlight correct and incorrect elements of student answers, as recognizing correct components can foster engagement and trust. To support error recognition, we will provide exemplary answers alongside LLM feedback to clarify why full credit was not earned and which concepts were misunderstood. CoTAL will be extended to identify misunderstandings for deeper discussion and to inform teachers so they can support struggling students. We will further strengthen stakeholder involvement through participatory design during CoTAL Phase II (Response Scoring, Prompt Development, and Active Learning).
7 Discussion
Our work provides several implications for K-12 educational methodology and practice. Conventional automated grading approaches, such as supervised fine-tuning, typically demand large annotated datasets and substantial GPU resources, rendering them costly and less adaptable to evolving classroom needs. In contrast, prompt-engineering methods like WRVRT [45] do not incorporate input from teachers or students and have not been evaluated comprehensively across multiple domains, limiting their pedagogical alignment and generalizability. CoTAL addresses these limitations by adopting a human-AI approach to aligning LLM behavior with teacher expectations and adapting responses to reflect the preferences of individual stakeholders in a generalizable manner. Crucially, the CoTAL approach achieves this using only a small set of demonstration instances to incorporate human insight, thereby reducing data and compute requirements while preserving cross-domain applicability. These properties position CoTAL as a practical, stakeholder-centered framework for formative assessment in K-12 contexts.
@@ 7 Discussion
This is particularly valuable for subjective tasks such as feedback generation and formative assessment scoring, where teachers’ preferences vary, making a single fine-tuned LLM unlikely to align well with multiple educators. While substantial research has focused on adapting LLMs to personalize student interactions, typically via content and feedback tailored to individual learning styles [62], CoTAL addresses a notable methodological gap by personalizing LLMs for educators. By incorporating teacher-specific preferences into scoring and feedback with minimal demonstration instances, CoTAL enables educator-aligned, context-sensitive evaluations that are difficult to achieve through conventional fine-tuning or generic prompting alone.
@@
Teacher and student feedback identified opportunities to strengthen stakeholder alignment and emphasized the necessity for cultivating trust in AI systems as a prerequisite for real-world adoption. Students frequently prefer ChatGPT-assisted (i.e., human-in-the-loop) grading over fully automated approaches [63, 64]. Consistent with this preference, Jiang et al. (2024) [60] reported general optimism about AI’s potential to enhance teaching among enduring concerns related to ethics, discrimination, and regulatory gaps. Lee and Zhai (2024) [65] further documented inappropriate use cases in which the LLM hallucinated internet material and cautioned that such use could “potentially undermine students’ practical inquiry skills over time.” These findings highlight the need for mechanisms that explicitly prioritize LLM reliability, trust, and curricular alignment—not merely accuracy—a conclusion reinforced by our results.
@@
Hallucination reduction is essential. Huang et al.’s (2023) [58] survey on LLM hallucinations identified several hallucination causes, discussing reasoning failure where “[an LLM] may struggle to produce accurate results if multiple associations exist between questions,” even in instances where the LLM possesses the necessary knowledge. In our case, this manifested as context inconsistency errors due to the LLM conflating the absorption and absorption limit concepts in the Rules and Debugging Tasks. These concepts are inextricably linked and often appear in similar contexts, which challenges differentiation. Just as students struggled to distinguish between these concepts, it is likely that the LLM’s training data reflects similar ambiguities, which we will investigate in future work. Hallucination mitigation strategies, such as retrieval-augmented generation (RAG) to ground responses in human-curated data and decoding strategies that prioritize adherence to facts and user instructions, similarly warrant future investigation.
We hypothesize that as foundation models and LLM methodologies continue to evolve, hallucinations will decline. Dhuliawala et al. (2024) [66] introduced Chain-of-Verification (CoVe), a reasoning-based approach in which the model generates verification questions from baseline responses to independently validate factual claims and produce corrected, evidence-aligned outputs. Their findings indicate that CoVe reduces hallucinations and improves alignment with task-relevant evidence. Recently, enterprise LLMs have adopted similar strategies, such as the reasoning capabilities integrated into OpenAI’s o3 and GPT-5 models.
However, despite the potential for advanced LLMs to reduce hallucinations, our recent work (discussed below) demonstrates that while reasoning-enabled models can help mitigate such errors, CoTAL remains an essential component of our automated scoring and feedback pipeline. Notably, factual hallucinations were significantly reduced—that is, LLMs adhered more closely to prompt constraints and appropriately cited evidence from students’ assessment responses. Nevertheless, formative assessment grading, evidence elicitation, and feedback generation are inherently subjective tasks. Even with LLM reasoning capabilities, multiple rounds of IRR evaluation, prompt refinement, and few-shot instance selection were necessary during active learning. This underscores the continued need for iterative and systematic prompt engineering for tasks where human agreement is challenging to establish.
Much of the misalignment between LLM predictions and human consensus in this work stemmed from overscoring (i.e., false positives). Given that LLMs are trained to please humans, it is unsurprising that they tend to award more points than warranted. However, in educational contexts, this behavior risks reinforcing misconceptions that may persist into future learning activities. OpenAI recently rolled back a GPT model update due to sycophantic behavior [67], highlighting the importance of examining this issue in CoTAL. In future work, we will extend our analysis to additional LLMs—including reasoning models and smaller, open-source alternatives— to better understand where our approach excels or falters across different families of models.
Our work is not without limitations. Although we demonstrated CoTAL’s generalizability across science, computing, and engineering within an integrated STEM+C curriculum, additional studies are necessary to evaluate its performance across broader tasks, domains, curricula, and student populations—although we note that our approach has since proven successful when applied to discourse analysis [41] and the nursing domain [68]. This study assessed CoTAL’s scoring performance in a post hoc context rather than in real-time classroom use, constraining our ability to infer its impact on student learning and classroom dynamics. Moreover, the human-in-the-loop prompt engineering required by CoTAL can be time-consuming—particularly during active learning—raising questions about scalability and sustainability. Future work must establish CoTAL’s capacity to reduce teacher workload and maintain alignment at scale in authentic classroom settings. Recent approaches, such as LLM-as-a-Judge [69] and LLM-as-a-Jury [70] show promise for evaluating LLMs towards this end.
@@
Recently, we conducted a follow-up study in four classrooms, each with approximately 26 students, deploying CoTAL via a formative assessment agent using refined formative assessments, rubrics, and prompts reflecting the student and teacher feedback gathered from this study (RQ2). This agent enabled students to engage with their CoTAL-generated formative assessment scores and explanations for the scores, helping them understand their mistakes and identify ways to improve their performance moving forward. While a comprehensive analysis of students’ learning gains and agent interactions is forthcoming, we (1) observed substantial time savings with CoTAL compared to our earlier human-scoring efforts—delivering feedback within hours instead of weeks, (2) leveraged this feedback to construct learner models based on evidence derived from the formative assessments [71], and (3) witnessed a considerable increase in LLM scoring performance across all assessments.
8 Conclusions